OpenStudy (anonymous):

x - y = 6 x = y + 2 Solve the system of equations by substitution. (4, -2) (6, 0) no solution coincident

OpenStudy (agent0smith):

You know what x is equal to, so plug that bad boy into the other equation.

OpenStudy (anonymous):

x = 8

OpenStudy (agent0smith):

@KarieMonsrerrxD you shouldn't give out answers, ESPECIALLY when they are wrong!

OpenStudy (anonymous):

lmao idc x would be 8 tho right ? because 8 - 2 is 6

OpenStudy (agent0smith):

Nope nope nope.

OpenStudy (anonymous):

oh , i tryed lol

jimthompson5910 (jim_thompson5910):

x - y = 6 (y+2) - y = 6 ... replace x with y+2 (since x = y+2) y+2 - y = 6 I'll let you finish

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

ohhhhhh

OpenStudy (anonymous):

hey magan

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

When you have x = y + 2 and you plug into x - y = 6 you get: (y + 2) - y = 6 ==> 2 = 6 (which can't be). If you do it the other way with y = x - 2 you would get: x - ( x - 2) = 6 ==> 2 = 6 (which also can't be). Therefore, there is no solution.